Solve initial value problem using Laplace transform

Problem 4 Solve the initial value problems given below --ез, y(0) 2. a. b. f ty 3 cos t, y(0)-
Answer #1

4a.

Given differential equation:

3t dt

Let's take Laplace transformation of both sides:

L\left ( \frac{dy}{dt} \right )-L(y)=L(e^{3t})

\Rightarrow \int _{0}^{\infty}dt \ \ e^{-st}\frac{dy}{dt}-L(y)=\int_{0}^{\infty}e^{(3-s)t}dt

\Rightarrow \int_{0}^{\infty}dt \frac{d}{dt}\left ( e^{-st}y \right )- \int_{0}^{\infty}y\frac{d}{dt}\left ( e^{-st} \right )dt -L(y)=\left[\frac{e^{(3-s)t}}{3-s} \right ]_0^\infty

\Rightarrow -y(0)+s \int_{0}^{\infty}y\left ( e^{-st} \right )dt -L(y)=-\left[\frac{1}{3-s} \right ]

\Rightarrow -2+sL(y)-L(y)=-\left[\frac{1}{3-s} \right ]

\Rightarrow -2+(s-1)L(y)=-\left[\frac{1}{3-s} \right ]

\Rightarrow (s-1)L(y)=-\left[\frac{1}{3-s} \right ]+2

\Rightarrow L(y)=-\left[\frac{1}{(3-s)(s-1)} \right ]+\frac{2}{s-1}

By inverse Laplace transformation,

y=2e^t-\frac{e^{3t}-e^t}{2}

--------------------------------------------------------------------------------------------------------------------------

3t dt